OpenStudy (needshelpdfdf):

Solve for x. x2 + 9x = 0

OpenStudy (eliesaab):

Factor \[ x^2+9x=x(x+9)=0 \]

OpenStudy (eliesaab):

Now it is easy

OpenStudy (needshelpdfdf):

0,9

OpenStudy (needshelpdfdf):

thank you can you help me with more

OpenStudy (eliesaab):

No 0 and -9

OpenStudy (needshelpdfdf):

yea that's what I meant